The probability that a blade manufactured by a factory is defective is 1/500. Blades are



packed in packets of 10 blades. Find the expected number of packets containing (i) no



defective blade (ii) one defective blade (iii) 2 defective blades, in a consignment of 10000



packets.



How can we solve this by using only poison distribution




1
Expert's answer
2022-04-07T08:41:35-0400

p="\frac{1}{500}" =0.002

per packet =0.002"\cdot" 10=0.02

Thus "\lambda" =0.02

P(X)="\frac{\lambda^xe^{-\lambda}}{x!}"

P(0)="\frac{0.02^0}{0!}\cdot e^{-0.02}=0.98"

The approximate # of packets with no defective blades is 0.98*10000=9800

P(1)="\frac{0.02^1}{1!}\cdot e^{-0.02}=0.0196"

The approximate # of packets with 1 defective blades is 0.0196*10000=196

P(2)="\frac{0.02^2}{2!}\cdot e^{-0.02}=0.0000196"

The approximate # of packets with 2 defective blades is 0.0000196*10000=2